OpenStudy (anonymous):

Find the exact circumference, in terms of , of a circle with the given radius. 7 feet

OpenStudy (anonymous):

\[14\pi \] 14.98

OpenStudy (anonymous):

sry 43.98

OpenStudy (bahrom7893):

2 pi * r = 2 * pi * 7 = 14pi so yea problemsolver is right.

OpenStudy (anonymous):

terry, wouldn't it be easier to just learn the formula, you have posted bout a 100 of these.

OpenStudy (bahrom7893):

good point too estudier..

OpenStudy (anonymous):

yea